# Rendering stable features improves sampling-based localisation with Neural radiance fields

### v1

@@ VI RESULTS AND DISCUSSION
Both Table I and the figures show that for ORB detection, rendering patches performs slightly better than rendering single pixels, aligning with iNeRF [15] findings that rendering interest regions increases performance.
MSER did not perform as well on the flower dataset, where random selection was more effective. This may be because the images in the flower dataset have many self-similar areas of red petals and green leaves (as shown in Fig. 1), which are less informative in stable regions. Random pixel selection is less vulnerable to this.
Table II shows the minimum number of network forward passes required to reach $\leq 10\%$ pose error for each dataset and method. The fern and flower datasets are quite difficult to localise in, as convergence in these complex real scenes with scattered features is hindered by the natural and amorphous objects there-in. Table entries without a number mean none of the runs of this feature selection method converged to under $\leq 10\%$ error. This is a relatively noisy set of measurements, as it does not consider aggregate performance over multiple runs (Table I is a more robust measure), but does give some indication of the orders of magnitude required for convergence. In general, with a sensible feature selection strategy, it is possible to reduce the number of network forward passes roughly tenfold.
TABLE II: Forward passes to achieve $\leq 10\%$ pose error
Dataset
@@
Note that most of the runs with few forward passes are from the runs that render 16 or 32 points along the ray, with around 30 initial poses on average. This indicates that when choosing between rendering high quality image pixels, for greater coverage of space, or smartly selecting the pixel to render, the latter plays a more important role for localisation performance.
VII LIMITATIONS AND FUTURE WORK
As mentioned above, the experiments conducted here are time consuming due to the large number of forward passes required for localisation. This limited our investigations to the pixel loss common in NeRF literature when constructing likelihoods. It may be that alternative strategies leveraging classical feature matching approaches or descriptors as patch likelihoods (eg. SSIM [30]) would be more effective, although these would require additional compute.
This work has focused on the inverse localisation problem and sampling-based localisation schemes. We hypothesise that alternative camera localisation strategies, for example training a pose estimation network directly using NeRF queries, are likely to equally effective. However, we believe that sampling-based approaches and the speed improvements here are still likely to be of significant value in robot planning settings relying on simulators or dynamics models and differentiable rendering pipelines [31].
VIII CONCLUSIONS
This paper investigated strategies to speed up sampling-based localisation with NeRFs. Our core finding is that rendering sampled stable features outperforms rendering approaches relying on corners or interest points and regions surrounding these. Results also show that reducing the number of points to be integrated along a NeRF ray does not significantly inhibit localisation, with rendered pixels and the number of poses sampled playing a more important role.

### v2

@@ 6.2 Results
Table 2: TE and RE of the experiment with 100 pixels and 45 poses
Translation Error (%) Rotation Error (%)
lego drums chair fern flower lego drums chair fern flower
Patch ORB 18.99 35.19 32.04 27.34 23.69 21.03 49.72 42.57 21.19 13.91
ORBrand 24.09 27.20 32.19 20.99 18.24 28.12 31.89 43.58 14.60 21.15
randfix 45.04 50.32 18.91 21.29 14.74 43.62 60.68 18.42 15.87 8.54
rand 6.22 19.45 14.42 23.12 22.83 4.95 15.83 12.68 4.19 5.18
MSER 4.37 17.41 4.18 21.47 10.84 2.94 16.29 3.56 6.76 10.23
MSERrand 1.28 9.25 13.16 14.29 22.38 1.10 7.31 17.02 2.57 19.98
Pixel ORB 22.94 38.43 29.64 22.37 21.20 25.48 57.36 37.80 13.10 13.51
ORBrand 33.42 30.70 32.12 25.66 16.02 41.89 37.02 42.34 16.53 17.92
randfix 4.60 5.68 15.77 42.01 12.80 3.07 4.72 14.93 4.97 10.56
rand 1.34 5.30 23.93 26.17 14.86 1.16 4.24 21.86 3.32 4.94
MSER 4.17 4.64 3.76 23.75 23.57 3.52 2.92 3.40 2.80 17.63
MSERrand 1.23 3.86 1.12 8.35 20.52 1.05 2.45 1.01 1.44 17.47
Figure 2 and Figure 3 show the final pose estimation error for each dataset after 20 iterations when rendering 16 and 64 points along the ray respectively, with little difference between these. [^3] Experiments with 32 points show similar patterns. In each heatmap, the x-axis shows varying numbers of initial pose particles with increasing pixel numbers rendered at each pose. The y-axis shows the different pixel sampling methods, with the same order as presented in the tables. Red lines separate experimental groupings for readability.
Similar performance holds when evaluating the translation and rotation errors separately. Table 2 presents the evaluation of one of the experiment settings when rendering 100 pixels and using 45 initial pose particles. This setting appears to provide the best balance between convergence and requiring a relatively low number of forward passes to complete. Results shows that pose estimation with MSER sampling method generally performs better across most datasets. When comparing with the performance of ORB sampling, the error percentage drops significantly. Similar trends are seen across other pixel and pose numbers.
@@
Both the tables and the figures show that for ORB detection, rendering patches performs slightly better than rendering single pixels, aligning with iNeRF [15] findings that rendering interest regions increases performance. This also supports our hypothesis since rendering (observing) a larger area around the corner points smooths the error likelihood.
MSER did not perform as well on the flower dataset, where random selection was more effective. This may be because the images in the flower dataset have many self-similar areas of red petals and green leaves (as shown in Figure 1), which are less informative in stable regions. Random pixel selection is less vulnerable to this.
In addition, Table 3 shows the minimum number of network forward passes required to reach $\leq 10\%$ pose error for each dataset and method. The fern and flower datasets are quite difficult to localise in, as convergence in these complex real scenes with scattered features is hindered by the natural and amorphous objects therein. Table entries without a number mean none of the runs of this feature selection method converged to under $\leq 10\%$ error. This is a relatively noisy set of measurements, as it does not consider aggregate performance over multiple runs (Table 1 is a more robust measure), but does give some indication of the orders of magnitude required for convergence. In general, with a sensible feature selection strategy, it is possible to reduce the number of network forward passes roughly tenfold.
Table 3: Forward passes to achieve $\leq 10\%$ pose error
@@
Note that none of the ORB selection method ever converged to below 10% estimation error. Most of the runs with few forward passes are from runs that render 16 or 32 points along the ray, with around 30 initial poses on average. This indicates that when choosing between rendering high quality image pixels, for greater coverage of space, or smartly selecting the pixel to render, the latter plays a more important role for localisation performance.
7 Limitations and Future Work
As mentioned above, the experiments conducted here are time consuming due to the large number of forward passes required for localisation. This limited our investigations to the pixel loss common in NeRF literature when constructing likelihoods. It may be that alternative strategies leveraging classical feature matching approaches or descriptors as patch likelihoods (eg. SSIM [35]) would be more effective, although these would require additional compute. Similarly, losses that make better use of patches (eg. SSIM) may be worth exploring, although rendering larger patches is likely to be too expensive for practical use. Our current experiments have also assumed static scenes and cameras. However, we expect that generalising to moving cameras would not change these results significantly.
8 Conclusion
This paper investigated strategies to speed up sampling-based localisation with NeRFs. Our core finding is that rendering sampled stable features, where image spatial changes are less frequent, outperforms rendering approaches relying on corners or interest points and regions surrounding these, where the feature changes rapidly. Results also show that reducing the number of points to be integrated along a NeRF ray does not significantly inhibit localisation, with rendered pixels and the number of poses sampled playing a more important role.